# Occupied Processes: Going with the Flow

### v3

@@ 7.1 Optimal Stopping
$v(o,x)=\sup_{\tau\in\vartheta(\operatorname{\mathcal{T}}_{t})}\operatorname{\mathbb{E}}^{\operatorname{\mathbb{Q}}}_{o,x}[\varphi(\operatorname{\mathcal{O}}_{\tau},X_{\tau})],\quad\operatorname{\mathcal{T}}_{t}=\operatorname{\mathcal{T}}\cap\ [t,T],\quad o(\operatorname{\mathbb{R}})=t,$ (90)
with $\operatorname{\mathbb{E}}^{\operatorname{\mathbb{Q}}}_{o,x}$ as in (36). The stopping region is traditionally defined as the coincidence set $\operatorname{\mathcal{S}}_{T}=\{(o,x)\in\operatorname{\mathcal{D}}_{T}\mathrel{\mathop{\ordinarycolon}}v(o,x)=\varphi(o,x)\}$. In particular, $\partial\operatorname{\mathcal{S}}=\partial\operatorname{\mathcal{D}}_{T}$ since $T\in\operatorname{\mathcal{T}}$. As it is challenging to represent $\operatorname{\mathcal{S}}$ either mentally or visually, we can regard the stopping region as a subset $\operatorname{\mathbb{R}} fo$r a given occupation measure, i.e., $\operatorname{\mathcal{S}}(o)=\{x\in\text{supp}(o)\ \mathrel{\mathop{\ordinarycolon}}\ (o,x)\in\operatorname{\mathcal{S}}\}$. For every $t\in[0,T]$ , this gives the pathwise partition
@@
with the continuation region $\operatorname{\mathcal{C}}(o)=\{x\in\text{supp}(o)\ \mathrel{\mathop{\ordinarycolon}}\ v(o,x)>\varphi(o,x)\}$. See Fig. 15 and [34, Chapter 7] for a similar comment.
The Dynamic Programming Principle (DPP) states that for every stopping time $\varsigma\in\vartheta(\operatorname{\mathcal{T}}_{t})$,
$v(o,x)$ $=\sup_{\tau\in\vartheta(\operatorname{\mathcal{T}}_{t})}\ \operatorname{\mathbb{E}}^{\operatorname{\mathbb{Q}}}_{o,x}[\varphi(\operatorname{\mathcal{O}}_{\tau},X_{\tau})\mathds{1}_{\{\tau<\varsigma\}}+v(\operatorname{\mathcal{O}}_{\varsigma},X_{\varsigma},)\mathds{1}_{\{\tau\geq\varsigma\}}].$ (91)
@@
$dY_{t}$ $=\Big{(}\partial_{o}+\frac{1}{2}\partial_{xx}\Big{)}v(\operatorname{\mathcal{O}}_{t},X_{t})dt+\partial_{x}v(\operatorname{\mathcal{O}}_{t},X_{t})dX_{t}.$ (92)
@@ 7.4 Shrinking the Corridor
Let $\varphi^{\varepsilon}(o,x)=\frac{o(B_{\varepsilon}(x))}{2\varepsilon}$, $\varepsilon>0$ as in (96). Hence $\varphi^{\varepsilon}(X_{t},\operatorname{\mathcal{O}}_{t})$ is a pathwise approximation of the spot local time $L_{t}^{X_{t}}$. We aim to show that the value $v_{*}^{\varepsilon}$ of the optimal stopping problem

### v4

@@ 7.1 Optimal Stopping
$v(\mathfrak{o},x)=\sup_{\tau\in\vartheta(\operatorname{\mathcal{T}}_{t})}\operatorname{\mathbb{E}}^{\operatorname{\mathbb{Q}}}_{\mathfrak{o},x}[\varphi(\operatorname{\mathcal{O}}_{\tau},X_{\tau})],\quad\operatorname{\mathcal{T}}_{t}=\operatorname{\mathcal{T}}\cap\ [t,T],\quad t=\mathfrak{o}(\operatorname{\mathbb{R}}).$ (90)
The stopping region is traditionally defined as the coincidence set $\operatorname{\mathcal{S}}_{T}=\{(\mathfrak{o},x)\in\operatorname{\mathcal{D}}_{T}\mathrel{\mathop{\ordinarycolon}}v(\mathfrak{o},x)=\varphi(\mathfrak{o},x)\}$. In particular, $\partial\operatorname{\mathcal{S}}=\partial\operatorname{\mathcal{D}}_{T}$ since $T\in\operatorname{\mathcal{T}}$. As it is challenging to represent $\operatorname{\mathcal{S}}$ either mentally or visually, we can regard the stopping region as a subset $\operatorname{\mathbb{R}} fo$r a given occupation measure, i.e., $\operatorname{\mathcal{S}}(\mathfrak{o})=\{x\in\text{supp}(\mathfrak{o})\ \mathrel{\mathop{\ordinarycolon}}\ (\mathfrak{o},x)\in\operatorname{\mathcal{S}}\}$. For every $t\in[0,T]$ , this gives the pathwise partition
@@
with the continuation region $\operatorname{\mathcal{C}}(\mathfrak{o})=\{x\in\text{supp}(\mathfrak{o})\ \mathrel{\mathop{\ordinarycolon}}\ v(\mathfrak{o},x)>\varphi(\mathfrak{o},x)\}$. See Fig. 15 and [34, Chapter 7] for a related comment.
The Dynamic Programming Principle (DPP) states that for all $\varsigma\in\vartheta(\operatorname{\mathcal{T}}_{t})$,
$v(\mathfrak{o},x)$ $=\sup_{\tau\in\vartheta(\operatorname{\mathcal{T}}_{t})}\ \operatorname{\mathbb{E}}^{\operatorname{\mathbb{Q}}}_{\mathfrak{o},x}[\varphi(\operatorname{\mathcal{O}}_{\tau},X_{\tau})\mathds{1}_{\{\tau<\varsigma\}}+v(\operatorname{\mathcal{O}}_{\varsigma},X_{\varsigma},)\mathds{1}_{\{\tau\geq\varsigma\}}].$ (91)
@@
$dY_{t}$ $=\Big{(}\partial_{\mathfrak{o}}+\frac{1}{2}\partial_{xx}\Big{)}v(\operatorname{\mathcal{O}}_{t},X_{t})dt+\partial_{x}v(\operatorname{\mathcal{O}}_{t},X_{t})dX_{t}.$ (92)
@@ 7.4 Shrinking the Corridor
Let $\varphi^{\varepsilon}(\mathfrak{o},x)=\frac{\mathfrak{o}(B_{\varepsilon}(x))}{2\varepsilon}$, $\varepsilon>0$ as in (96). Hence $\varphi^{\varepsilon}(X_{t},\operatorname{\mathcal{O}}_{t})$ is a pathwise approximation of the spot local time $L_{t}^{X_{t}}$. We aim to show that the value $v_{*}^{\varepsilon}$ of the optimal stopping problem